OpenStudy (anonymous):

the right triangle whose hypotenuse is sqrt6 m long is revolved about one of its legs to generate a right circular cone. find the radius, height, and the volume of the cone of the greatest volume that can be made.

OpenStudy (phi):

sounds like calculus

OpenStudy (phi):

with height h and radius r being the legs of a right triangle, you have h^2 + r^2= 6 Volume of a cone is V= (1/3) pi r^2 h sub in r^2= (6-h^2) and take the derivative wrt h Set to zero and solve for h